Shift Summary for day, Nov. 29, 1999

Angela Bellavance (bellavan@rice.edu)
Mon, 29 Nov 1999 15:36:16 -0600 (CST)

Shift Summary for day shift, 29/Nov/99

On Shift: L. Bellantoni, A. Bellavance

Hours of beam delivered at all intensities: 0
Hours of beam delivered at nominal intensity: 0
Hours of E799 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: TeV still recovering from power glitch

========================================================================
Runs ended this shift: several, but inconsequential

Tapes brought to FCC this shift: UPA205-7, UPB205-7, UPC205-7, UPS069,
UPE011-4, UPH026-35,UPF038-51,UPPE06-7,
UPCR03, UPLA08, UPK020-6, UPM031-40,
UPX014-5, UPT012-2

========================================================================
Accesses during shift: Valeri and Sasha L. to work on VV' latches.
Sydney and Jason H. to work on ET.

========================================================================
Overview of shift: waiting for beam, debugging VV' latches, and chasing
gremlins... lots of gremlins...

========================================================================
Description of Shift:

time comment
------------------------------------------------------------------------
09:12 Dave N. notified about gas flow. He says don't worry about it.
09:40 Beam est. pushed back to 2pm, so incomplete muon run stopped and
cosmics started for more debugging by Sasha L. and Valeri.
10:54 Bryon alarm goes off - panic ensues. Turns out to be Leonard
testing the alarm system without telling us.
12:00 Peter Shanahan remotely installs new laser timing.
12:23 Etotal tripped off. Sydney and Jason H. called in to look at
it.
13:45 All the broken junk on the table is taken away by a nice man
from Prep named Kevin.
14:00 Ch. 13: "Start up and turn on activities commencing." MCR calls
to ask what mode we want to come up in (muon). Might we get
beam soon?
ET reset from LV trip, value tweeked. Noticed temperature on
channel 9 thermometer is 76 degrees (6 degrees too high).
Leonard paged. Claims there is no problem but will have the set
points for the hall temp lowered 3 degrees.
14:35 Interlocks inexplicably drop. Called MCR - sending techs.
14:43 TRD alarms all go off at once. Fortunately Sasha L. is right
here.
15:03 People arrive to investigate interlocks.
15:25 The interlocks were not dropped. The circuit breaker which
powers the indicator light for the interlocks was tripped.
Circuit reset. This trip was also responsible for the TRD
alarms and the ET alarms.

========================================================================
Follow up Notes on Previous Shift

- muon run _not_ done yet (no beam yet).
- junk taken to prep.
- tapes vaulted.
- keys found in parking lot taken to Wilson Communications Center (lost
and found).

========================================================================
Items to Follow up on Next Shift

- do muon run.
- don't worry about the Ar-Eth flowmeters being low unless they drop below
50.

_______
Angela Bellavance work: (630) 840-4144
bellavan@rice.edu home: (630) 548-3412
http://www.owlnet.rice.edu/~bellavan/ page: (630) 722-6861
"It has yet to be proven that intelligence has any survival value."
- Arthur C. Clarke